A compound H2X with molar weight of 80g is dissolved in a solvent having density of 0.4g ml^-1 Assume no change in volume upon dissolution, the molality of a 3.2 molar solution is

Grade:11

1 Answers

Ritesh Kumar
6 years ago
Molarity = moles of solute/vol.of soln----1Molality =moles of solute/wt of solvent --21/2>M/m=wt of solvent/vol. Of solnA/c vol does not change so, M/m = density3.2/m= 0.4m=8(m)